y=4x-9
y=-x+6
Solve a system by substitution

Answer by checkley75(3666)

Y=4X-9 NOW SUBSTITUTE(4X-9) FOR Y IN THE OTHER EQUATION & SOLVE FOR X.
4X-9=-X+6
4X+X=6+9
5X=15
X=15/5
X=3 ANSWER.
Y=4*3-9
Y=12-9
Y=3 ANSWER.
PROOF
3=-3+6
3=3